OpenStudy (anonymous):

The perimeter of a rectangle is 148 m. The length is 2 m more than twice the width. Find the dimensions. The length is ?

OpenStudy (anonymous):

2l + 2w = 148 2(2w + 2) + 2w = 148 4w + 4 + 2w = 148 6w + 4 = 148 6w = 144 w = 144/6 w = ?

OpenStudy (anonymous):

w=24 and l is (148-2(24))/2=50